OpenStudy (anonymous):

Abel cycled a total distance of 2 and 1 over 2 miles during the months of March and April. If Abel only cycled 1 over 8 of a mile every day, which expression shows the number of days he cycled? 2 and 1 over 2 ⋅ 1 over 8 2 and 1 over 2 + 1 over 8 2 and 1 over 2 – 1 over 8 2 and 1 over 2 ÷ 1 over 8

OpenStudy (anonymous):

@Ibbutibbu.

OpenStudy (anonymous):

@IMStuck

OpenStudy (anonymous):

@DangerousJesse

OpenStudy (imstuck):

Hold on a sec...I'll be right with you, ok?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@IMStuck

OpenStudy (anonymous):

u there m8

OpenStudy (anonymous):

@DangerousJesse

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

@e.mccormick

OpenStudy (texaschic101):

(2 1/2) / (1/8) = (5/2) / (1/8) = 5/2 * 8/1 = 40/2, which reduces to 20 days check : 1/8 * 20 = 20/8 = 2 1/2 ...yep, thats it :)

OpenStudy (anonymous):

so which one is it

OpenStudy (anonymous):

i think its a

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

@IMStuck

OpenStudy (anonymous):

@texaschic101